A Bibliography of Publications in Supercomputing Review

Nelson H. F. Beebe Center for Scientific Computing

University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org, beebe@ieee.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

02 December 2011 Version 1.03

Title word cross-reference

1 [Ano92-35].

3 [Ano92u].

’92 [Ano92f].

Accelerate [Ano92l]. Acceptance [Can91b]. Across [Smi92f]. Add [Smi91b]. Add-On [Smi91b]. Affected [All91a]. After [Ano92u]. Again [Smi91a]. Agree [Ano92d]. Agrees [Ano92k]. Ahead [Ano92a]. AIDS [Eis92]. Alamos [Ano92v]. Algorithm [BB91]. Always [Gre92a].

America [Ami91]. Analysis [Bur92a]. Ancient [All91a]. Animation [Ano91b]. Apples [Smi91d]. Applications [Row91]. Applied [Ano92a]. Approaches [Pat92]. Arsenide [Hil91]. Art [Sma91a]. Asteroids [Smi92a]. Astronomy [BS91]. Atomic [Pat92]. Atomic-Scale [Pat92]. Attempts [Can92c]. Awaits [Ano92z]. Awards [Ano92h].

Back [Ano91b]. Barcelona [Ano91g]. Based [Ruw91]. Basics [Ano92-29]. Be [Ano91b]. Beach [Smi91c]. Beat [Can92c]. Benchmark [GREC90]. Benchmarking [BB91]. Benchmarks [Cyb91]. Better [BB91]. Beyond [Pat92]. Big [Sma91d].

Billion [Ano91f]. Biomedicine [Row91]. Blue [Ano91c]. Board [Smi92f]. Bottleneck [LB92a]. Breaking [BS91]. Breakthrough [Pat91]. Bright [Bur91c]. Bring [Law92]. Brink [Sto91]. Broader [Nev92]. Built [Ano91g]. Bush [Ano92b]. Buyer [Ano91a].

C90 [Ano92u]. CAE [Can91a]. Can [Ano92j]. Cancels [Ano92u]. Catalyst [Can91b]. CDC [Ano92d]. Center [Ano92-32]. Centers [Sch91a]. Championships [Rat91a]. Charting [MTYYs92]. Cheap [Rat92d]. Checklist [Nil91]. Cheers [Ano91c]. Chess [Rat91a]. Chief [Ano92t]. Chip [Bur92d]. Chipmaker [Hil91]. Chooses [Ano91j]. Claims [Ano92x]. Claus [Sma92b]. Climate [Ano91i]. Climates [All91a]. Commercial [Tet91]. Community [Ano92q]. Comparison [Sim91]. Competition [Bur91c]. Compiler [LB92a]. Computational [Her92]. Compute [Her91d]. Compute-Intensive [Her91d]. Computer [Ano91e, Ano92u]. Computers [Bai91]. Computing [Her91a]. Conceptual [Sma91a]. Conference [Ano92-27]. Conferenced [Sma91b]. Connectivity [Smi92b]. Conquer [Rat91d]. Consortia [Ano92i]. Consortium [Ano92m]. Context [Nev92]. Contract [Ano92-32]. Contraire [Kal92]. Control [Ano92d]. Convex [Ano92e]. Corning [Tet91]. Corp. [Ano92x]. Corporate [Ami91]. Cost [MA91]. Cost-Effective [Ruw91]. Couple [Sch91b]. Cray [Ano92g]. Cray-3 [Ano92u]. CSPP [Ano92l]. CTP [Ano92w]. Curriculum [Ano92y].

D [Her92]. DARPA [Ano92i].

Data [Bur92a, Ano92d]. Days [Ano91g]. DEC [Ano92-34]. December [Ano92u]. Deep [Rat91a]. Delta [Ano92j]. Demo [Ano92u]. Demonstrates [Ano92j]. Design [Bur92d]. Deskside [Smi91b]. Despite [Bur91c]. Develop [Ano92s]. Developing [Bur92d]. Devices [SFJ91a]. Diamond [Law92]. Digital [Ano92k]. Discusses [Ano92e]. Disk [Ruw91]. Disk-Based [Ruw91]. Dispels [Hil91]. Divide [Rat91d]. DOE [Ano92v]. Doing [Smi92d]. Door [BDG^{+}91]. Drip [BS92b]. Drop [BS92b]. Dug [Rat91a]. Dump [Bur92a]. Duration [Rat91a]. Dynamics [Her92].

Earth [MTYYs92]. Easy [Rat92d].

Economies [Ano91k]. Educating [Bur92e]. Education [Bur92e]. Edward [Ano92t]. Effective [MA91]. Effort [Ano92h]. Encounters [Can91a]. End [Bur91a]. Enhance [Ive92a]. Enters [Ano91h]. Environmental [Pat91]. Environments [Ive92a]. Epidemiology [Eis92]. Equations [Bur92c]. Europe [Ano91f]. European [Ano92m]. Evaluation [Sim91]. Evolution [Tet91]. Exclusive [Her92]. Expectations [Sch91a]. Experience [Ano92j]. Expertise [VC92a]. Exploration [Law91].

Face [Smi91c]. Factory [Bur92d]. Fall [Sch91a]. Fermilab [Her91d]. Financiers [Bur92a]. First [GREC90]. Five [Sim91]. Fluid [Her92]. Focus [Can92a]. Fool [Bai91]. Forces [Ano92s]. Forecast [Ano91f]. FORGEs [Ano92a]. Form [Ano92i]. FPS [Smi91a]. France [Smi92d]. Fray [Ano92-34]. Freeze [Ano92-31]. Future [Bur91c]. FY’93 [Ano92b].

GaAs [Hil91]. Gallium [Hil91]. Games [Sma92a]. Geared [Ano92y].

Generation [Bur92e]. Geoscience [Pat91]. Giving [Bai91]. Going [Smi92f]. Gold [Smi92f]. Good [Rat92d]. Good-Looking [Rat92d]. Graphics [Ano91f, Ano91h]. Great [Sch91a]. Green [Ano91c]. Grow [Ano92-30]. Guessing [Sma92a]. Guide [Ano91a].

Handy [Smi92b]. Hardware [Ano92c]. Heterogeneous [Smi91a]. High [Ano91j]. High-End [Bur91a]. High-Performance [LB92a]. High-Speed [Ano91j]. Highly [Bai91]. Hire [VC92a]. Hitachi [Ano92n]. HP [Ano92-33]. HPC [Ano92i]. HPCC [Ano92b]. Hungry [Smi91b]. Hypertasking [Bab91].

i860 [Ano92o]. IBM [Ano91j]. Iceberg [Ano92-31]. If [FP91]. II [Fur91]. I’ll [Ano91b]. Impact [Bur92b]. Increase [Ano91k]. Indigo [Rat91b]. Industrial [Sch91a]. Industry [Sch91a]. Information [Ano92q]. Innovations [Law92]. Insights [Sma91c]. Integrated [Ano91h]. Integrators [VC92a]. Intel [Ano92h]. Intensive [Her91d]. Interactive [Ruw91]. Interior [MTYYs92]. International [Ano91k]. Interview [Ano92t]. Isn’t [Sma92b]. Ivory [Rat92a]. Ivory-Tower [Rat92a].

Japan [Her91a]. Japanese [Fur91]. Join [Ano92g]. Joins [Ano92p].

Kalb [Gre92b]. Keeping [Smi92c].

Labs [Ano92v]. Lags [Ano92u]. LawnMower [Bur91d]. Leader [Can91a]. Leading [Ano91k]. Licensing [Grz91]. Life [All91a]. Livermore [Ano92u]. Looking [Rat92d]. Low [Ano92-34]. Low-Cost [Ano92-34]. Low-End [Ano92k].

Machine [Ano92m]. MacNeal [Can91a]. MacNeal-Schwendler [Can91a]. Making [BS91]. Man [Bur91d]. Market [Ano91h]. Marketplace [Bur91b]. Markets [Ano91f]. Masi [Ano92t]. Mass [Ano92-28]. Masses [Bai91]. Massively [Nil91]. Math [Sto91]. Mathematics [Sto91]. Maximum [Ano91j]. Means [Can92a]. Meant [Ano91b]. Medicine [Row91]. Meets [Her91d]. Meiko [Ano92w]. Melt [Ano92-31]. Memory [LB92a]. Microchip [Pat92]. Microprocessors [Smi92a]. Microsystems [Can92b]. Midst [Ano92p]. Million [Ano92h]. MIPS [Ano91e]. Modeling [Eis92]. Moves [Ano92i]. MP [Ano92k]. Mpp [Her91d, Can91b]. Multiprocessor [Smi91d]. Muscle [Smi91c]. Mutations [Rat91e]. Myths [Hil91].

Named [Ano92v]. NCSA [Ive92b]. NEC [Ano92d]. Need [Her91d]. Network [Kat91]. Networking [Ano92-29]. Networks [Kat91]. Neural [Kat91]. Neurocomputer [Kat91]. Never [Smi91h]. Neves [Gre92b]. Newport [Ano92-27]. News [Ano91e]. Next [Bur92e]. No [Sma92b]. Not-So-Odd [Sch91b]. Numerically [Tet91].

Oak [Ano92v]. Odd [Sch91b]. Odds [Can92c]. Olympics [Smi92f]. Once [Smi91a]. Opening [BDG^{+}91]. Opens [Grz91]. Operability [Smi92b]. Opts [Ano92u]. Orange [Ano91c]. Order [Ano92u]. OSF [Ano92-35]. OSF/1 [Ano92-35]. OTHERWARE [Rat92a]. Outlook [Ano91f]. Outranks [Ano92-33].

Pacesetter [Can92b]. Paracompiler [Bab91]. Parade [Ano92p]. Parallel [Bai91]. Parsytec [Ano92z].

Part [Fur91]. Particle [Her91d]. Partnership [Ano92f]. Partnerships [Sch91a]. Peak [Ano92x]. 6000 [Ano91j]. Percent [Ano91k]. Perfect [Cyb91]. Performance [Cyb91]. Perspective [Bai91]. Perspectives [BB91]. Pervades [Ano92-27]. Pessimism [Ano92-27]. Petroleum [Tch92]. Picture [Sma91d]. Pioneers [Smi91a]. Pittsburgh [Sch91b]. Place [Rat92b]. Planets [Smi92a]. Plans [Ano92m]. Player [Ano91e]. Pollution [Pat91]. Pontifications [Rat92a]. Possibility [Kat91]. Power [Smi91b]. Power-Hungry [Smi91b]. Predicted [Ano92-30]. Preemptive [Sma91c]. Price [Smi92e]. Prime [Kal92]. PRINTWARE [Rat92a]. Problems [Ano92w]. Program [Ano92b]. Programs [Ano92y]. Project [Ano92p]. Projections [Sch91a]. Projects [Bur91c]. Prudential [Row92]. Putting [Bur92c].

R [Her92]. RAID [Ano91j]. Raises [Ano92p]. Ready [Kal92]. Real [Rat92a]. Real-World [Rat92a]. Reality [Bur91d]. Reconstruct [All91b]. Red [Ano91c]. References [Smi92b]. Refining [VC92b]. Relationship [VC92b]. Release [Ano92f]. Releases [Bur91a]. Remodeled [Ano92x]. Reorganization [Ano92p]. Reply [Gre92b]. Requests [Ano92b]. Research [Ano92v, Ano92g]. Researcher [Bur92d]. Resource [Ano92q]. Results [Bai91]. Ridge [Ano92v]. RS [Ano91j]. RS/6000 [Ano91j]. Russian [Tch92].

S [Can92a]. Said [Ano91b]. Sale [Ano92w]. Sales [Ano91k]. Santa [Sma92b]. Scalable [GREC90]. Scale [Pat92]. Scenes [Bur91d]. Schwendler [Can91a]. Scientific [Ruw91]. SCInet [SR91].

Sculptural [Rat91e]. Sector [Tet91]. Securities [Row92]. Seeing [Row91]. Seeking [BS92a]. Seeks [Ano92f]. Self [Her91d]. Self-Wrought [Her91d]. Sell [Ano92k]. Semiconductor [SFJ91a]. Sequences [Ano91b]. Servers [Smi91d]. Sets [Smi91h]. SGI [Bur91a]. Shape [Row91]. Shelving [Ano92f]. Ships [Ano92-36]. Short [Sch91a]. Showcase [Ano91h]. Showing [Kat91]. Signs [Ano92-32]. Silicon [Ano91h]. SIMD [Smi91b]. Simulations [SFJ91a]. Simulator [Ano91h]. Skywriter [Ano91h]. SLALOM [GREC90]. Small [Ano91e]. Smaller [Rat92b]. Softgraph [Smi91f]. Software [Bab91]. Solution [Her91d]. Solutions [Ruw91]. Sony [Smi91c]. Space [Law91]. SPARC [Ano92g]. SPARC-Based [Ano92g]. SPARCbook [Ano92-36]. Special [Ano92-28]. Specialized [Gre92a]. Speed [Ano91j]. Speeds [Ano92x]. Spotlights [Ano92w]. Spread [Eis92]. Spreadsheets [Bur92a]. Standard [Ano92w]. Standards [Ano92-29]. Stanford [Bur92d]. Stardent [Smi91g]. Stars [BS91]. Status [Ano92p]. Storage [Ano91j, Ano92-31]. Strategies [Ano92j]. Strategy [Ano91j]. Street [Row92]. Stress [Ano92o]. Strikes [Sma91c]. Study [Ano92-33]. Submicron [SFJ91a]. Success [Nil91]. Sudden [Bur92b]. Sun [Bur91c]. Super [Kat91]. SuperCompilers [Ano92o]. Supercomputer [GREC90, Sch91a]. Supercomputers [All91b]. Supers [Rat91c]. Superserver [Ano92g]. Supplement [Ano92-28]. Sweeping [Ano92p]. SX [Ano92d]. SX-3 [Ano92d]. System [Ano91d]. Systems [Ano91e].

Table [Ano91e]. Tadpole [Ano92-36]. Take [Pat91]. Target [Law91].

Techniques [Ive92b]. Technologies [Ano92-36]. Technology [Smi91a]. Tek [Ano92-31]. Ten [Smi92e]. Tera [Can92c]. Teraflops [Ano92h]. Terminator [Ano91b]. Theory [Her91d]. There [Sma92b]. Things [Row91]. Thinking [Smi92c]. Thought [Rat91a]. Three [Ano91c]. Time [Kal92]. Tools [LB92a]. Top [Ano91e]. Total [Ano91f]. Tower [Rat92a]. Transfer [Ano92l]. Trends [Cyb91]. Tries [Smi91g]. Troubled [Can91a]. Twelve [Bai91].

U.S. [Ano91f]. Uncertain [Smi91a]. Unix [Ano92-30]. US$10.6 [Ano92-32]. US$21 [Ano92h]. US$210 [Ano91h]. US$43 [Ano91f]. USA [Smi92d].

Vectorization [Ano92o]. Vendors [Fur91]. Venus [Law91]. Violet [Ano91c]. Virginia [Sma92b]. Virtual [Bur91d]. Visual [Bur92a]. Visualization [Ano91d]. Visualizations [Rat91c].

Wall [Row92]. Wars [Bur91b]. was [Ano91g]. Waters [Can91a]. Wavetracer [Smi91b]. Ways [Bai91]. Weather [Ano91i]. Well [Smi92d]. Whole [Smi91e]. Will [Ano92-31]. Wired [SR91]. Wobbly [Ano91k]. Worchester [Ano92-32]. Word [Smi91g]. Work [Ano92j]. Workstation [Ano91f]. Workstations [Ano91f]. World [Rat92a]. Worth [Ano92-32]. Wrought [Her91d].

Y-MP [Ano92k]. Years [Smi92e]. Yellow [Ano91c].

Zephyr [Smi91b].

References

Allen:1991:ACH

[All91a] William H. Allen. Ancient climates: How ancient climates affected life. Supercomputing Review, 4(6):20–??, June 1991. CO-DEN SURVEG. ISSN 1048-6836.

Allen:1991:ACS

[All91b] William H. Allen. Ancient climates: Supercomputers reconstruct. Supercomputing Review, 4(6):20–??, June 1991. CODEN SURVEG. ISSN 1048-6836.

Amini:1991:CPS

[Ami91] Mohammad M. Amini. Corporate perspective: Supercomputing in corporate America. Supercomputing Review, 4(9):40– 42, September 1991. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1991:BGW

[Ano91a] Anonymous. Buyers’ guide: Workstations and visualization. Supercomputing Review, 4(10): 48–??, October 1991. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1991:SSI

[Ano91b] Anonymous. Showcase: He said, “I’ll Be Back,” and meant it: Terminator II computer animation sequences. Supercomputing Review, 4(9):25–??, September 1991. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1991:STC

[Ano91c] Anonymous. Showcase: Three cheers for the red, orange, yellow, green, blue and violet. Supercomputing Review, 4(6):59–??, June 1991. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1991:WNI

[Ano91d] Anonymous. Workstation news:

IBM’s new visualization system. Supercomputing Review, 4(8):46– ??, August 1991. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1991:WNM

[Ano91e] Anonymous. Workstation news:

MIPS computer systems: Small player at the top table. Supercomputing Review, 4(6):43–??, June 1991. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1991:WOB

[Ano91f] Anonymous. Workstation outlook: US$43 billion total U.S. and Europe markets for graphics workstations forecast for 1995. Supercomputing Review, 4(6): 42–??, June 1991. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1991:WSB

[Ano91g] Anonymous. Workstation showcase: Barcelona was built in 28 days. Supercomputing Review, 4 (9):48–??, September 1991. CO-DEN SURVEG. ISSN 1048-6836.

Anonymous:1991:WSS

[Ano91h] Anonymous. Workstation showcase: Silicon Graphics enters simulator market with US$210,000 integrated Skywriter. Supercomputing Review, 4(6):48–??, June 1991. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1991:WSW

[Ano91i] Anonymous. Workstation showcase: Weather and climate. Supercomputing Review, 4(8):48–??, August 1991. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1991:WSI

[Ano91j] Anonymous. Workstation storage: IBM chooses maximum strategy for RS/6000 RAID highspeed storage. Supercomputing Review, 4(6):44–??, June 1991. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1991:WWS

[Ano91k] Anonymous. The workstation word: Sales by leading workstation vendors increase 40 percent despite wobbly international economies. Supercomputing Review, 4(10):44–??, October 1991. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:APR

[Ano92a] Anonymous. Applied parallel research FORGEs ahead. Supercomputing Review, 5(5):13–??, May 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:BRP

[Ano92b] Anonymous. Bush requests 23 percent increase for HPCC program in FY’93. Supercomputing Review, 5(3):10–??, March 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:BGH

[Ano92c] Anonymous. Buyers’ guide: Hardware and systems. Supercomputing Review, 5(1):56– ??, January 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:CDN

[Ano92d] Anonymous. Control Data and NEC agree on CDC sales of SX-3 supercomputers. Supercomputing Review, 5(2):10–??, February 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:CDM

[Ano92e] Anonymous. Convex discusses MPP plans. Supercomputing Review, 5(5):11–??, May 1992. CO-DEN SURVEG. ISSN 1048-6836.

Anonymous:1992:CCS

[Ano92f] Anonymous. Cray Computer shelving Cray-3 release in ’92; seeks partnership. Supercomputing Review, 5(3):8–??, March 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:CRS

[Ano92g] Anonymous. Cray Research and Sun to join in SPARC-based superserver project. Supercomputing Review, 5(2):8–??, February 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:DAM

[Ano92h] Anonymous. DARPA awards US$21 million to Intel’s teraflops effort. Supercomputing Review, 5 (3):7–??, March 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:DMF

[Ano92i] Anonymous. DARPA moves to form HPC consortia. Supercomputing Review, 5(2):14–??, February 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:DED

[Ano92j] Anonymous. Delta experience demonstrates MPP and HPCC strategies can work. Supercomputing Review, 5(5):7–??, May 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:DAS

[Ano92k] Anonymous. Digital agrees to sell low-end CRAY Y-MP EL. Supercomputing Review, 5(2):14– ??, February 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:DCA

[Ano92l] Anonymous. DOE and CSPP accelerate technology transfer. Supercomputing Review, 5(5):14–??, May 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:ECP

[Ano92m] Anonymous. European consortium plans to release teraflops machine in 1993. Supercomputing Review, 5(3):13–??, March 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:HCT

[Ano92n] Anonymous. Hitachi claims top speeds for new supercomputers. Supercomputing Review, 5 (5):12–??, May 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:ISS

[Ano92o] Anonymous. i860 supercompilers stress vectorization. Supercomputing Review, 5(5):16–??, May 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:MSR

[Ano92p] Anonymous. In midst of sweeping reorganization, IBM raises status of supercomputing, joins MPP parade with new project. Supercomputing Review, 5(2):6– 7, February 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:IRHa

[Ano92q] Anonymous. The information resource for the high-performance computing community. Supercomputing Review, 5(4):57, April 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:IRHb

[Ano92r] Anonymous. The information resource for the high-performance computing community. Supercomputing Review, 5(5):57–??, May 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:IDJ

[Ano92s] Anonymous. Intel and Digital join forces to develop MPP software. Supercomputing Review, 5(5):8–??, May 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:IEM

[Ano92t] Anonymous. Interview: Edward Masi, Intel’s new supercomputer chief. Supercomputing Review, 5 (5):32–35, May 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:LCC

[Ano92u] Anonymous. Livermore cancels Cray-3 order; opts for C90 after Cray Computer lags on December demo. Supercomputing Review, 5(2):9–??, February 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:AOR

[Ano92v] Anonymous. Los Alamos and Oak Ridge Labs named DOE Research Centers. Supercomputing Review, 5(2):12–??, February 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:MSS

[Ano92w] Anonymous. Meiko sale spotlights problems with CTP standard. Supercomputing Review, 5 (2):17–??, February 1992. CO-DEN SURVEG. ISSN 1048-6836. Anonymous:1992:NCC

[Ano92x] Anonymous. NEC Corp. claims top peak speeds for remodeled SX-3. Supercomputing Review, 5 (3):11–??, March 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:NEP

[Ano92y] Anonymous. New education programs geared toward the curriculum. Supercomputing Review, 5 (2):16–??, February 1992. CO-DEN SURVEG. ISSN 1048-6836. Anonymous:1992:PMA

[Ano92z] Anonymous. Parsytec MPP awaits chip. Supercomputing Review, 5(5):16–??, May 1992. CO-DEN SURVEG. ISSN 1048-6836. Anonymous:1992:PPN

[Ano92-27] Anonymous. Pessimism pervades Newport conference. Supercomputing Review, 5(5):10–??, May 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:SSM

[Ano92-28] Anonymous. Special supplement:

Mass storage. Supercomputing Review, 5(3):42–??, March 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:SSN

[Ano92-29] Anonymous. Special supplement:

Networking: Basics and standards. Supercomputing Review, 5(5):41–??, May 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:UEP

[Ano92-30] Anonymous. Unix in Europe predicted to grow. Supercomputing Review, 5(5):15–??, May 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:WST

[Ano92-31] Anonymous. Will Storage Tek’s new Iceberg freeze IBM, or melt first? Supercomputing Review, 5 (3):12–??, March 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:WCS

[Ano92-32] Anonymous. Worchester center signs MPP contract worth US$10.6 million. Supercomputing Review, 5(3):14–??, March 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:WCH

[Ano92-33] Anonymous. Workstation comparison: HP outranks competition in workstation study. Supercomputing Review, 5(1):49– ??, January 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:WNDa

[Ano92-34] Anonymous. Workstation news: DEC enters low-cost workstation fray. Supercomputing Review, 5 (1):46–??, January 1992. CO-DEN SURVEG. ISSN 1048-6836.

Anonymous:1992:WNDb

[Ano92-35] Anonymous. Workstation news: DEC releases industry’s first OSF/1 OS. Supercomputing Review, 5(3):40–??, March 1992. CODEN SURVEG. ISSN 1048-6836.

Anonymous:1992:WNT

[Ano92-36] Anonymous. Workstation news: Tadpole Technologies ships SPAR-Cbook. Supercomputing Review, 5(3):40–??, March 1992. CODEN SURVEG. ISSN 1048-6836.

Baber:1991:PSS

[Bab91]  Marc Baber.  Parallel soft-

ware solutions: the hypertasking paracompiler. Supercomputing Review, 4(11):41–??, November 1991. CODEN SURVEG. ISSN 1048-6836.

Bailey:1991:HPP

[Bai91] David H. Bailey. Highly parallel perspective: Twelve ways to fool the masses when giving performance results on parallel computers. Supercomputing Review, 4(8):54–??, August 1991. CO-DEN SURVEG. ISSN 1048-6836.

Bjorstad:1991:BPS

[BB91] Petter E. Bjorstad and Erik Boman. Benchmarking perspectives: Slalom: A better algorithm. Supercomputing Review, 4 (11):57–??, November 1991. CO-DEN SURVEG. ISSN 1048-6836.

Beguelin:1991:WSO

[BDG^{+}91] Adam Beguelin, Jack Dongarra, Al Geist, Robert Manchek, and Vaidy Sunderam. Workstation solutions: Opening the door to heterogeneous network supercomputing. Supercomputing Review, 4(9):44–45, September 1991. CODEN SURVEG. ISSN 1048-6836.

Barnes-Svarney:1991:SAM

[BS91] Patricia Barnes-Svarney. Supercomputing in astronomy: The making and breaking of stars. Supercomputing Review, 4(9):28– 32, September 1991. CODEN SURVEG. ISSN 1048-6836.

Barnes-Svarney:1992:SS

[BS92a] Patricia Barnes-Svarney. Seeking the Sun. Supercomputing Review, 5(4):20–??, April 1992. CODEN SURVEG. ISSN 1048-6836.

Barnes-Svarney:1992:WSW

[BS92b] Patricia Barnes-Svarney. Workstation showcase: When a drip is a drop. Supercomputing Review, 5(1):54–??, January 1992. CO-DEN SURVEG. ISSN 1048-6836.

Burns:1991:SRH

[Bur91a] Mathew Burns. SGI releases high-end graphics system. Supercomputing Review, 4(12):42– ??, December 1991. CODEN SURVEG. ISSN 1048-6836.

Burns:1991:SMS

[Bur91b] Mathew Burns. The supercomputing marketplace: Scientific visualization wars. Supercomputing Review, 4(10):31–35, October 1991. CODEN SURVEG. ISSN 1048-6836.

Burns:1991:WNS

[Bur91c] Mathew Burns. Workstation news: Sun projects bright future despite competition. Supercomputing Review, 4(9):43– ??, September 1991. CODEN SURVEG. ISSN 1048-6836.

Burns:1991:WSV

[Bur91d] Mathew Burns. Workstation showcase: Virtual virtual reality: Scenes from the LawnMower Man. Supercomputing Review, 4 (11):48–??, November 1991. CO-DEN SURVEG. ISSN 1048-6836.

Burns:1992:FDS

[Bur92a] Mathew Burns. Financiers dump spreadsheets for visual data analysis. Supercomputing Review, 5 (4):27–??, April 1992. CODEN SURVEG. ISSN 1048-6836.

Burns:1992:NSS

[Bur92b] Mathew Burns. News showcase: Sudden impact. Supercomputing Review, 5(2):22–??, February 1992. CODEN SURVEG. ISSN 1048-6836.

Burns:1992:PEW

[Bur92c] Mathew Burns. Putting equations to work. Supercomputing Review, 5(5):26–??, May 1992. CODEN SURVEG. ISSN 1048-6836.

Burns:1992:SRD

[Bur92d] Mathew Burns. Stanford researcher developing “virtual factory” for chip design. Supercomputing Review, 5(4):42–??, April 1992. CODEN SURVEG. ISSN 1048-6836.

Burns:1992:SEE

[Bur92e] Mathew Burns. Supercomputers in education: Educating the next generation. Supercomputing Review, 5(1):39–??, January 1992. CODEN SURVEG. ISSN 1048-6836.

Canin:1991:OIM

[Can91a] Jeffrey Canin. Outside insights: MacNeal-Schwendler: CAE leader encounters troubled waters. Supercomputing Review, 4(11):24–??, November 1991. CODEN SURVEG. ISSN 1048-6836.

Canin:1991:OIC

[Can91b] Jeffry Canin. Outside insights:

Catalyst for MPP market acceptance. Supercomputing Review, 4 (9):22–??, September 1991. CO-DEN SURVEG. ISSN 1048-6836.

Canin:1992:OIF

[Can92a] Jeffrey Canin. Outside insights:

Focus means success for E and S. Supercomputing Review, 5(1): 26–??, January 1992. CODEN SURVEG. ISSN 1048-6836.

Canin:1992:OIS

[Can92b] Jeffry Canin. Outside insights:

Sun Microsystems: Workstation pacesetter. Supercomputing Review, 5(3):16–??, March 1992.

CODEN SURVEG. ISSN 1048-6836.

Canin:1992:TCA

[Can92c] Jeffry Canin. Tera Computer attempts to beat the odds. Supercomputing Review, 5(5):18–??, May 1992. CODEN SURVEG. ISSN 1048-6836.

Cybenko:1991:SPT

[Cyb91] George Cybenko. Supercomputer performance trends and the Perfect Benchmarks. Supercomputing Review, 4(4):53–60, April 1991. CODEN SURVEG. ISSN 1048-6836.

Eisenberg:1992:SEM

[Eis92] Steve Eisenberg. Supercomputers in epidemiology: Modeling the spread of AIDS. Supercomputing Review, 5(3):34–??, March 1992. CODEN SURVEG. ISSN 1048-6836.

Freund:1991:NSI

[FP91] Richard F. Freund and Larry J. Peterson. Network solutions: If the network is the computer, then. Supercomputing Review, 4 (10):36–??, October 1991. CO-DEN SURVEG. ISSN 1048-6836.

Furutsuki:1991:JSJ

[Fur91] Tad Furutsuki. Japanese supercomputing: The Japanese supercomputing market, part II: The Japanese vendors. Supercomputing Review, 4(8):42–45, August 1991. CODEN SURVEG. ISSN 1048-6836.

Gregory:1992:MPW

[Gre92a] Peter Gregory. MPP perspective: Will MPP always be specialized? Supercomputing Review, 5(3):28–??, March 1992. CODEN SURVEG. ISSN 1048-6836.

Gregory:1992:RKN

[Gre92b] Peter Gregory. A reply to Kalb and Neves. Supercomputing Review, 5(5):62–??, May 1992. CO-DEN SURVEG. ISSN 1048-6836.

Gustafson:1990:SFS

[GREC90] John Gustafson, Diane Rover, Stephen Elbert, and Michael Carter. SLALOM—the first scalable supercomputer benchmark. Supercomputing Review, 3(11): 56–61, November 1990. CODEN SURVEG. ISSN 1048-6836.

Grzanka:1991:WMS

[Grz91] Len Grzanka. Workstation marketplace: Silicon Graphics opens GL licensing. Supercomputing Review, 4(11):50–??, November 1991. CODEN SURVEG. ISSN 1048-6836.

Herbst:1991:JSPs

[Her91a] Kris Herbst. Japanese supercomputing: Parallel computing in Japan, part I. Supercomputing Review, 4(9):33–35, September 1991. CODEN SURVEG. ISSN 1048-6836.

Herbst:1991:JSPa

[Her91b] Kris Herbst. Japanese supercomputing: Parallel computing in Japan, part I. Supercomputing Review, 4(9):33–??, September 1991. CODEN SURVEG. ISSN 1048-6836.

Herbst:1991:JSPb

[Her91c] Kris Herbst. Japanese supercomputing: Parallel computing in Japan, part II. Supercomputing Review, 4(10):38–39, October 1991. CODEN SURVEG. ISSN 1048-6836.

Herbst:1991:SPT

[Her91d] Kris Herbst. Supercomputing in particle theory: Fermilab’s self-wrought mpp solution meets compute-intensive need. Supercomputing Review, 4(8):58–64, August 1991. CODEN SURVEG. ISSN 1048-6836.

Herbst:1992:JRE

[Her92] Kris Herbst. Japanese R and D Exclusive: Computational fluid dynamics in Japan. Supercomputing Review, 5(1):30–??, January 1992. CODEN SURVEG. ISSN 1048-6836.

Hill:1991:GAC

[Hil91] Richard H. Hill. Gallium arsenide chipmaker dispels GaAs myths. Supercomputing Review, 4(6):26–??, June 1991. CODEN SURVEG. ISSN 1048-6836.

Iverson:1992:SVE

[Ive92a] Wes Iverson. Showcase: Virtual environments enhance. Supercomputing Review, 5(1):28– ??, January 1992. CODEN SURVEG. ISSN 1048-6836.

Iverson:1992:SVT

[Ive92b] Wes Iverson. Showcase: Visualization techniques at NCSA. Supercomputing Review, 5(1):28– ??, January 1992. CODEN SURVEG. ISSN 1048-6836.

Kalb:1992:MPM

[Kal92] Jeffrey C. Kalb. MPP perspective: MPP specialized? au contraire: It’s ready for prime time. Supercomputing Review, 5 (4):48–??, April 1992. CODEN SURVEG. ISSN 1048-6836.

Kato:1991:NNN

[Kat91]  Hideki Kato.  Neural net-

works: Neural network simulations showing the possibility of a super neurocomputer. Supercomputing Review, 4(6):38–??, June 1991. CODEN SURVEG. ISSN 1048-6836.

Lawton:1991:SET

[Law91] George Lawton. Space exploration: Target: Venus. Supercomputing Review, 4(12):40– ??, December 1991. CODEN SURVEG. ISSN 1048-6836.

Lawton:1992:DIB

[Law92] George Lawton. Diamond innovations bring new power to computers. Supercomputing Review, 5(4):38–??, April 1992. CODEN SURVEG. ISSN 1048-6836.

Loshin:1992:CTHa

[LB92a] David Loshin and Dan Budge.

Compiler tools in high-performance computing: Breaking the memory bottleneck, part 1. Supercomputing Review, 5(1):43–??, January 1992. CODEN SURVEG. ISSN 1048-6836.

Loshin:1992:CTHb

[LB92b] David Loshin and Dan Budge. Compiler tools in high-performance computing: Breaking the memory bottleneck, part 2. Supercomputing Review, 5(2):56–??, February 1992. CODEN SURVEG. ISSN 1048-6836.

Menasce:1991:HSC

[MA91] Daniel Menasce and Virgilio Almeida. Heterogeneous supercomputing: Cost effective? Supercomputing Review, 4(8):39–41, August 1991. CODEN SURVEG. ISSN 1048-6836.

Morin:1992:CEI

[MTYYs92] Paul J. Morin, Toshiro Tanimoto, David A. Yuen, and Zhang Yushen. Charting the Earth’s interior. Supercomputing Review, 5 (2):36–??, February 1992. CO-DEN SURVEG. ISSN 1048-6836.

Neves:1992:BC

[Nev92] Kenneth W. Neves. A broader context. Supercomputing Review, 5(5):58–??, May 1992. CODEN SURVEG. ISSN 1048-6836.

Nilsson:1991:MPS

[Nil91] Bob Nilsson. MPP perspectives: a software checklist for massively parallel success. Supercomputing Review, 4(12):50–??, December 1991. CODEN SURVEG. ISSN 1048-6836.

Patterson:1991:BGS

[Pat91] Kent Patterson. Breakthrough in geoscience: Supercomputers take on environmental pollution. Supercomputing Review, 4(11):35– ??, November 1991. CODEN SURVEG. ISSN 1048-6836.

Patterson:1992:BMT

[Pat92] Kent Patterson. Beyond the microchip: Three approaches to atomic-scale memory devices. Supercomputing Review, 5(4):30–??, April 1992. CODEN SURVEG. ISSN 1048-6836.

Rathbone:1991:CCC

[Rat91a] R. Andrew Rathbone. Computer chess championships: Deep Thought II dug in for the duration. Supercomputing Review, 4 (12):6–??, December 1991. CO-DEN SURVEG. ISSN 1048-6836.

Rathbone:1991:NSS

[Rat91b] Tina Rathbone. News showcase:

Silicon Graphics’ new Indigo system. Supercomputing Review, 4 (9):20–??, September 1991. CO-DEN SURVEG. ISSN 1048-6836.

Rathbone:1991:NSV

[Rat91c] Tina Rathbone. News showcase:

Visualizations on new supers. Supercomputing Review, 4(12):35– ??, December 1991. CODEN SURVEG. ISSN 1048-6836.

Rathbone:1991:SDC

[Rat91d] Tina Rathbone. Showcase: Divide and conquer. Supercomputing Review, 4(11):26–??, November 1991. CODEN SURVEG. ISSN 1048-6836.

Rathbone:1991:SSM

[Rat91e] Tina Rathbone. Showcase: Sculptural mutations. Supercomputing Review, 4(10):24–??, October 1991. CODEN SURVEG. ISSN 1048-6836.

Rathbone:1992:POV

[Rat92a] Tina Rathbone. PRINTWARE and OTHERWARE: Virtual reality: Real-world applications and ivory-tower pontifications. Supercomputing Review, 5(1):66– ??, January 1992. CODEN SURVEG. ISSN 1048-6836.

Rathbone:1992:SMW

[Rat92b] Tina Rathbone. Showcase: Making the world a smaller place. Supercomputing Review, 5(3):24–??, March 1992. CODEN SURVEG. ISSN 1048-6836.

Rathbone:1992:SV

[Rat92c] Tina Rathbone. The software vendors. Supercomputing Review, 5(5):21–??, May 1992. CODEN SURVEG. ISSN 1048-6836.

Rathbone:1992:WSC

[Rat92d] Tina Rathbone. Workstation showcase: Cheap, easy and goodlooking. Supercomputing Review, 5(2):52–??, February 1992. CO-DEN SURVEG. ISSN 1048-6836.

Rowell:1991:AMP

[Row91] Jan Rowell. Applications in medicine: Parallel supercomputers: Seeing the shape of things in biomedicine. Supercomputing Review, 4(11):28–??, November 1991. CODEN SURVEG. ISSN 1048-6836.

Rowell:1992:PSP

[Row92] Jan Rowell. Prudential Securities pioneers supercomputing on Wall Street. Supercomputing Review, 5(4):24–26, April 1992. CODEN SURVEG. ISSN 1048-6836.

Ruwart:1991:CSH

[Ruw91]  Thomas M. Ruwart.  Cost-

effective solutions: High-speed, disk-based interactive scientific animation. Supercomputing Review, 4(9):36–??, September 1991. CODEN SURVEG. ISSN 1048-6836.

Schatz:1991:SIG

[Sch91a] Willie Schatz. Supercomputing and industry: Great expectations: Industrial partnerships fall short of Supercomputer Centers’ projections. Supercomputing Review, 4(11):65–68, November 1991. CODEN SURVEG. ISSN 1048-6836.

Schneider:1991:PNC

[Sch91b] Michael Schneider. Pittsburgh’s not-so-odd couple. Supercomputing Review, 4(8):36–??, August 1991. CODEN SURVEG. ISSN 1048-6836.

Shur:1991:SSSa

[SFJ91a] Michael Shur, Tor A. Fjeldy, and Geir U. Jensen. Supercomputer simulations of submicron semiconductor devices. Supercomputing Review, 4(6):27–32, June 1991. CODEN SURVEG. ISSN 1048-6836.

Shur:1991:SSS

[SFJ91b] Michael Shur, Tor A. Fjeldy, and Geir U. Jensen. Supercomputer simulations of submicron semiconductor devices. Supercomputing Review, 4(6):27–??, June 1991. CODEN SURVEG. ISSN 1048-6836.

Simmons:1991:WCP

[Sim91] Margaret L. Simmons. Workstation comparison: a performance evaluation of five workstations. Supercomputing Review, 4 (10):40–??, October 1991. CO-DEN SURVEG. ISSN 1048-6836.

Smaby:1991:OICb

[Sma91a] Gary Smaby. Outside insights: Conceptual art. Supercomputing Review, 4(12):38–??, December 1991. CODEN SURVEG. ISSN 1048-6836.

Smaby:1991:OICa

[Sma91b] Gary Smaby. Outside insights: Conferenced out. Supercomputing Review, 4(8):22–??, August 1991. CODEN SURVEG. ISSN 1048-6836.

Smaby:1991:OIP

[Sma91c] Gary Smaby. Outside insights: Preemptive strikes. Supercomputing Review, 4(6):18–??, June 1991. CODEN SURVEG. ISSN 1048-6836.

Smaby:1991:OIB

[Sma91d] Gary Smaby. Outside insights: the big picture. Supercomputing Review, 4(10):22–??, October 1991. CODEN SURVEG. ISSN 1048-6836.

Smaby:1992:GG

[Sma92a] Gary Smaby. Guessing games.

Supercomputing Review, 5(4): 22–??, April 1992. CODEN SURVEG. ISSN 1048-6836.

Smaby:1992:OIN

[Sma92b] Gary Smaby. Outside insights:

No, Virginia, there isn’t a Santa Claus. Supercomputing Review, 5(2):24–??, February 1992. CO-DEN SURVEG. ISSN 1048-6836.

Smith:1991:HSO

[Smi91a] Norris Parker Smith. Heterogeneous supercomputing: Once again, FPS pioneers uncertain new technology. Supercomputing Review, 4(6):33–37, June 1991. CODEN SURVEG. ISSN 1048-6836.

Smith:1991:WZS

[Smi91b] Norris Parker Smith. Wavetracer’s Zephyr: A SIMD deskside add-on for the power-hungry. Supercomputing Review, 4(12): 42–??, December 1991. CODEN SURVEG. ISSN 1048-6836.

Smith:1991:WNS

[Smi91c] Norris Parker Smith. Workstation news: Sony’s new MPP face on muscle beach. Supercomputing Review, 4(10):43–??, October 1991. CODEN SURVEG. ISSN 1048-6836.

Smith:1991:WWB

[Smi91d] Norris Parker Smith. The workstation word: Blue apples and multiprocessor servers. Supercomputing Review, 4(8):52–??, August 1991. CODEN SURVEG. ISSN 1048-6836.

Smith:1991:WWSd

[Smi91e] Norris Parker Smith. The workstation word: Seeing things whole. Supercomputing Review, 4 (12):44–??, December 1991. CO-DEN SURVEG. ISSN 1048-6836.

Smith:1991:WWSb

[Smi91f] Norris Parker Smith. The workstation word: Softgraph. Supercomputing Review, 4(9):46– ??, September 1991. CODEN SURVEG. ISSN 1048-6836.

Smith:1991:WWSa

[Smi91g] Norris Parker Smith. The workstation word: Stardent tries again. Supercomputing Review, 4 (6):46–??, June 1991. CODEN SURVEG. ISSN 1048-6836.

Smith:1991:WWSc

[Smi91h] Norris Parker Smith. The workstation word: Sun never sets: New servers and software. Supercomputing Review, 4(11):55– ??, November 1991. CODEN SURVEG. ISSN 1048-6836.

Smith:1992:MPA

[Smi92a] Norris Parker Smith. Microprocessors: Planets and asteroids. Supercomputing Review, 5 (5):39–??, May 1992. CODEN SURVEG. ISSN 1048-6836.

Smith:1992:POC

[Smi92b] Norris Parker Smith. PRINT-WARE and OTHERWARE: Connectivity and operability: Handy references. Supercomputing Review, 5(2):58–??, February 1992. CODEN SURVEG. ISSN 1048-6836.

Smith:1992:POK

[Smi92c] Norris Parker Smith. PRINT-WARE and OTHERWARE: Keeping up with mass storage and thinking about MPP. Supercomputing Review, 5(4):50–??, April 1992. CODEN SURVEG. ISSN 1048-6836.

Smith:1992:WWI

[Smi92d] Norris Parker Smith. The workstation word: IBM: Doing well in the USA and in France. Supercomputing Review, 5(3):41–??, March 1992. CODEN SURVEG. ISSN 1048-6836.

Smith:1992:WWT

[Smi92e] Norris Parker Smith. The workstation word: Ten years of Sun: Price, price, software, software. Supercomputing Review, 5 (4):47–??, April 1992. CODEN SURVEG. ISSN 1048-6836.

Smith:1992:WWWb

[Smi92f] Norris Parker Smith. The workstation word: Workstation Olympics: HP going for the gold across the board. Supercomputing Review, 5(2):50–??, February 1992. CODEN SURVEG. ISSN 1048-6836.

Smith:1992:WWWa

[Smi92g] Norris Parker Smith. The workstation word: Workstations and supercomputers. Supercomputing Review, 5(1):50–??, January 1992. CODEN SURVEG. ISSN 1048-6836.

Smith:1991:SIW

[SR91] Norris Parker Smith and Tina Rathbone. SCInet had it wired. Supercomputing Review, 4(12):2– ??, December 1991. CODEN SURVEG. ISSN 1048-6836.

Storad:1991:SMM

[Sto91] Conrad J. Storad. Supercomputing in mathematics: Math on the brink. Supercomputing Review, 4 (8):26–31, August 1991. CODEN SURVEG. ISSN 1048-6836.

Tchashchin:1992:HCR

[Tch92]  Kirill Tchashchin.  High-

performance computing in the Russian petroleum industry. Supercomputing Review, 5(2):32– ??, February 1992. CODEN SURVEG. ISSN 1048-6836.

Teter:1991:SCS

[Tet91] Michael Teter. Supercomputing in the commercial sector: the evolution of numerically intensive computing at Corning. Supercomputing Review, 4(10):27– 30, October 1991. CODEN SURVEG. ISSN 1048-6836.

Vivoli:1992:SIPa

[VC92a] Martha Vivoli and Dana Chamberlain. Systems integrators, part I: Expertise for hire. Supercomputing Review, 5(2):27–??, February 1992. CODEN SURVEG. ISSN 1048-6836.

Vivoli:1992:SIPb

[VC92b] Martha Vivoli and Dana Chamberlain. Systems integrators, part II: Refining the relationship. Supercomputing Review, 5(3):19–??, March 1992. CODEN SURVEG. ISSN 1048-6836.